# Attention-based Class-Conditioned Alignment for Multi-Source Domain Adaptive Object Detection

### v1

@@ 3 Proposed Method
where $d_{K}$ is the size of the key $K$. The dot-product of $K$ and $Q$ allows our model to learn the relation between the object features and class labels based on object category and image domain. Later, the dot-product of this with $V$, allows our model to learn which part of the object feature to focus on for class-conditioned adaptation. We also combine multiple single-head attention to get multi-head attention. This allows our method to learn multiple such relations. At the end of the attention block, a classification head is used to predict the domain of the input instance. The instance aligner is trained with the cross-entropy loss:
$\mathcal{L}_{\mbox{cdis}}=-\sum_{j=1}^{N}\sum_{i=1}^{M_{i}}\frac{1}{Z_{i}^{j}}\sum_{r=1}^{Z_{i}^{j}}d_{i}^{j}\log(D(R(F(z_{i,r}^{j}))))\ ,$ (5)
where $D$, $R$, and $F$ represent the Discriminator, ROI-Pooling, and Feature extractor of the student model; $Z_{i}^{j}$ is the number of Ground Truth(GT) instances in $i$th image from domain $j$, and $z_{i,r}^{j}$ is the $r$th GT box on $i$th image from domain $j$. The target of the discriminator, $d_{i}\in\{0,1\}^{N}$, is the domain label for each GT box as in Eq. 3. Note that $d_{i}$ is of length $N$ because we are not considering (pseudo) GTs from the target image.
Final Training Objective: After starting the student-teacher mutual learning all the losses discussed above are combined and used for training the student model. So, the final training loss for the student model would be:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyper-parameters for weighing each losses. The teacher model is updated as an EMA of the student model at each iteration.
4 Results and Discussion
@@ 4.1 Experimental Methodology
1. Implementation Details - In our experiments, we follow the same settings as [35, 32, 3]. In all experiments, Faster-RCNN[22] with Imagenet pretrained VGG-16 backbone is used as the detector. Images are resized such that the shorter side of the image is 600 while maintaining the image ratios, following the ROI-alignment[12] based implementation of Faster-RCNN. For the hyperparameters, we set $\alpha$, $\beta$ and $\gamma$ to 1, 0.1 and 0.3, respectively. The pseudo-label threshold was set to 0.7 for all the experiments. As mentioned in 3, our training is two-staged. We trained our model for 15 epochs in both the settings. We set the weight smoothing coefficient for EMA of teacher mode to 0.9996. All experiments were done on 4 A100 GPUs, with batch size and learning rate of 8 and 0.2 respectively. We implemented our method using Pytorch.
@@
2. Baseline Methods - We compared our results against four settings: 1 - Lower Bound - Supervised training on the merged source data. 2 - UDA - Source domains are blended and the UDA method is used. 3 - MSDA - SOTA MSDA methods fall into this category. 4 - Upper Bound - We have 3 upper bounds: Target Only - Supervised training on target data, All-Combined - Supervise training on target and source combined, Fine-Tuning - Trained on source domain, then fine-tuned on the target domain.
@@
3. Datasets and Adaptation Setting For our experiments we use 5 different datasets with various objects and different domains and capturing conditions: BDD100K [36], Cityscapes [9], KITTI [11], MS COCO [18] and Synscapes [31]. The MSDA problem settings used in the empirical study are summarized in Tab. 1. For each setting, the source and target domains used are listed as well.
Setting Src.1 # Img. Src.2 # Img. Src.3 # Img. Target # Img.
Cross-time Day 36,728 Night 27,971 - - Dawn 5,027
Cross-camera Cityscapes 2,831 Kitty 6,684 - - Day 36,728
Table 1: Summary of the different MSDA settings used in our study.
4.2 Results
1. Cross-time Adaptation - Daytime and Night subsets of BDD100K dataset are used as source domain, while the Dusk/Dawn subset of BDD100K is used as target domain. In this setting the images are captured at different times. This causes a domain shift due to variation in illumination.
@@ 4.2 Results
We reported the result of our method, against other baseline in Tab. 2. It can be observed that the DAOD methods outperforms the Lower Bound. But, there is a slight decrease in performance when the Night domain is added to the Daytime domain. This is because the UDA Blending methods did not consider the domain shift among the source domains, resulting in a degradation of performance. This emphasizes the importance of research on the MSDA methods. Our method beats the previous state of art MSDA method[3] by 2.6 mAP. Also, our method outperforms Target-Only and All-Combined Upper Bound Settings. The poor performance of the Target Only setting is due to fewer samples in the target domain compared to source domains.
@@
2. Cross-camera Adaptation- Cityscapes and Kitty datasets are used as source domains, while the Daytime subset of BDD100K is used as the target domain. In this setting the images are captured using different cameras. This causes a domain shift due to variations in the camera’s resolution and viewpoint. In this setting, the training and evaluation is done only on the car object category (the only overlapping class across the domains).
The AP on car compared against the baselines is reported in Tab. 2. The general variation of performance in this setting is similar to Sec. 4.2 setting. Our proposed method is able to outperform the previous method by 0.4%. Our result is also very close to the target-only upper-bound. In this setting there is only one object category, thus the class-conditioning has no impact in instance alignment. As shown in the ablation studies, the class-conditioning during instance alignment is crucial for the performance improvement.
Setting Method D N D+N C K C+K
Lower Bound Source Only 30.4 25.0 28.9 44.6 28.6 43.2
UDA Strong-Weak[26] 31.4 26.9 29.9 45.5 29.6 41.9
Graph Prototype[4] 31.8 27.6 30.6 - - -
Cat. Regularization[33] 31.2 28.4 30.2 46.5 30.8 43.6
UMT[10] 33.8 21.6 33.5 47.5 35.4 47.0
Adaptive Teacher[16] 34.9 27.8 34.6 49.8 40.1 48.4
MSDA MDAN[38] - - 27.6 - - 43.2
[…]
DMSN[35] - - 35.0 - - 49.2
TRKP[32] - - 39.8 - - 58.4
PMT[3] - - 45.3 - - 58.7
ACIA (Ours) - - 47.9 - - 59.1
Upper Bound Target-Only - - 26.6 - - 60.2
All-Combined - - 45.6 - - 69.7
Fine-Tuning - - 50.9 - - 72.1
@@
Table 2: Detection mAP of our method compared against the baselines in the cross-time (left) and cross-camera (right) setting. Source domains are daytime (D) and night (N) subsets and the target is Dusk/Dawn subset of BDD100K in cross-time setting. And, for the cross-camera setting source domains are Cityscapes (C) and Kitty (K) and the target is Daytime subset of BDD100K.
3. Extension to Mixed Domain Adaptation - MS COCO, Cityscapes, and Synscapes datasets are used as source domain. While, the Daytime subset of BDD100K is used as target domain. In the real-world data the domain shift is not limited to a single factor. So, in this setting domain shift occurs due to multiple and variable reasons. Also, in this setting, we have three source domains, instead of two. These reasons make this setting challenging and very close to practical setting.
@@
The empirical results from this setting is presented in Tab. 3. In this setting the Lower Bound is performing better than the UDA Blending baseline. We hypothesize this is due to large domain shift among domains so blending all the source domain as one and using UDA is resulting in performance degradation. This highlights the difficulty of this setting. When compared with the previous MSDA methods, our method is outperforming them. When Cityscapes and MS COCO are used as source domain, we’re outperforming the previous SOTA result by 2.3 mAP. When Synscapes is also added to the source domain this number rises to 2.6 mAP. This shows that our method continues to improve the performing when source domain are added. Our method is also outperforming the Target Only Upper Bound.
Setting Method C C+M C+M+S
@@
Table 3: mAP on 7 object categories of our method compared against baselines in the mixed setting. Source domains are Cityscapes(C), MS COCO(M), and Synscapes(S) datasets while the Daytime domain of BDD100K is target domain.
@@ 4.3 Ablation Studies
Unless mentioned, all the ablation studies are performed on the cross-time adaptation setting as in [35, 3]. The impact of the components performing image-level and instance-level feature alignment is analyzed along with other possible design choices for class-conditioned instance alignment.
1. Impact of the adaptation components - First, we analyzed the impact of the individual components of our proposed method. These components include image-level discriminator and instance-level discriminator. To better understand the importance of class conditioning in the instance alignment, we further examined the instance classifier with and without class-wise attention. The results are reported in Tab. 4. It can be observed that just by adding the image-level discriminator, the performance improved by a huge margin. Furthermore, by adding the instance-level discriminator (without class-embedding) the performance of the model increased by 3.4% showing the importance of instance-level alignment. The performance is further boosted by integrating the class-information into the instance-level alignment.
@@
w/o class-embed. w/ class-embed.
@@
Table 4: Impact of individual components our ACIA method. Image-LA: Image-Level Alignment, Instance-LA: Instance-Level Alignment.
2. Design choices for class-wise instance alignment - The instance-level alignment component can be designed in different ways. Apart from the attention-based method, class information can be incorporated to instances by concatenation and multiplication. Tab. 5(left) reports the variation in performance with different design choice. It is clear that attention is the superior design choice. We believe this is because we’re learning just one class-embedding layer across every domain. When we’re merging the class-information to the ROI-Pool features, attention can learn what information is more important for a particular domain in a better way than multiplication and concatenation. This is because of the design of the attention layer, allowing it to learn complex relations [27].
@@
Table 5: (left) Different design choices for integrating class information in instance alignment; (right) Detection performance degrades when target domain is used in the instance alignment process.
3. Impact of target domain on class-wise alignment - Instance-level alignment differs from image-level alignment in that it focuses solely on the foreground region of the image. In this section, we investigate the impact of employing or omitting the target image in instance-level alignment. As reported in Tab. 5 (right), utilizing target data in both scenarios leads to a decline in detector performance. In a DA scenario, the target data lack labels, making it challenging to distinguish between foreground regions and object categories. One approach to address this challenge is by employing pseudo-labels; however, they often introduce noise and degrade performance. Nonetheless, in MSDA, the availability of multiple source domains enables us to choose which domains to align at the instance level. Hence, our work focuses on aligning solely the source domains.
4. Detection Visualization In Fig. 4 we present the visualization of detections on BDD100k cross-time for the three multi-source adaptation approaches presented in the paper: no class-wise adaptation, prototype-based class-conditional adaptation and our attention-based class-conditional adaptation.
5. Level of Imbalance In Tab. 6 we compare our method against PMT for different levels of imbalance. As expected, when the dataset is balanced the two methods perform similarly. However, when the level of imbalance increases, our method obtains much better mAP. This shows that the error accumulation from the pseudo-labels is more problematic in the case of imbalanced datasets. Our class-conditioned attention model solves that issue, as it uses the same model for all classes and therefore can better deal with high levels of imbalance.
@@
(a) No class-wise (b) Prototype based (c) Ours
Figure 4: Examples of ODs on BDD100k cross-time setting, with different types of instance-level adaptation. (a) ODs with a multi-source adaptation without instance level class-conditional adaptation as in [35, 32]. (b) ODs with the prototype-based class-conditional adaptation [3]. (c) ODs with our ACAI approach.
Method Bike Bus Car Motor Person Rider Light Sign Truck mAP
Mostly Balanced 25.0 50.0 10.0 15.0
PMT 43.8 52.3 - 28.0 - 34.4 - - - 39.6
Ours 45.8 48.9 - 30.7 - 31.0 - - - 39.1
Slightly Imbalanced 4.6 9.2 1.8 59.6 2.8 22.0
PMT 42.1 52.7 - 25.1 43.1 34.0 - - 50.5 41.2
Ours 44.3 54.2 - 25.3 45.3 40.6 - - 50.4 43.4
Mostly Imbalanced 1.2 2.4 0.5 15.4 0.7 32.8 42.3 5.7
PMT 49.2 56.2 - 29.6 49.9 32.0 43.2 55.8 55.3 46.4
Ours 52.1 59.2 - 29.2 53.1 36.1 46.2 58.1 57.1 48.9
Highly Imbalanced 0.5 1 57.8 0.2 6.5 0.3 13.4 17.8 2.4
PMT 55.3 59.8 67.6 29.9 47.6 32.7 46.3 56.0 57.7 50.3
Ours 56.1 61.0 69.2 31.9 51.8 39.8 49.2 59.0 61.0 53.2
Table 6: Detection results per-class on cross-time setting for PMT and our method with different levels of imbalance. The dataset is quite imbalanced. To create a more balanced dataset we reduced the number of classes by removing those that are creating the imbalance. For each level of imbalance, we report the relative importance of each class in percentage.

### v2

@@ 3 Proposed Method
where $d_{K}$ is the size of the key $K$. The dot-product of $K$ and $Q$ allows our model to learn the relation between the object features and class labels based on object category and image domain. Later, the dot-product of this with $V$, allows our model to learn which part of the object feature to focus on for class-conditioned adaptation. The multi-head attention allows our method to learn multiple such relations. At the end of the attention block, a classification head is used to predict the domain of the input instance. The instance aligner is trained with the cross-entropy loss:
$\mathcal{L}_{\mbox{cdis}}=-\sum_{j=1}^{N}\sum_{i=1}^{M_{i}}\frac{1}{Z_{i}^{j}}\sum_{r=1}^{Z_{i}^{j}}d_{i}^{j}\log(D_{I}(R(F(z_{i,r}^{j}))))\ ,$ (5)
where $D_{I}$, $R$, and $F$ represent the instance-level discriminator, ROI-pooling, and feature extractor of the student model, $Z_{i}^{j}$ is the number of Ground Truth (GT) instances in the $i$th image from domain $j$, and $z_{i,r}^{j}$ is the $r$th GT box on the $i$th image from domain $j$. $d_{i}^{j}$ is a one-hot vector for each GT box as in Eq. 3. Note that $d_{i}^{j}$ has length $N$ because we do not consider (pseudo) GTs from the target images.
(e) Overall Training Objective. In the student-teacher mutual learning stage, all the losses discussed above are combined, and the overall training loss for the student model is:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyper-parameters for weighting losses. The teacher is updated as an EMA of the student at each iteration as $w_{t}=\delta w_{t}+(1-\delta)w_{s}$ where $w_{t},w_{s}$ are the weights of teacher and student models respectively, and $\delta$ is a smoothing factor.
4 Experiments
@@ 4.1 Experimental Methodology
(a) Implementation Details. In our experiments, we follow the same settings as [36, 33, 3]. In all experiments, Faster-RCNN [23] is used as the detector, with Imagenet pre-trained VGG-16 backbone. Images are resized such that the shorter side of the image is 600 while maintaining the image ratios, following the ROI-alignment [13] based implementation of Faster-RCNN. We set $\alpha$, $\beta$, and $\gamma$ hyperparameters to 1, 0.1, and 0.3, respectively. The pseudo-label threshold was set to 0.7 for all the experiments. As mentioned in Sec. 3, our training is two-staged. We trained our model for 15 epochs in both the settings. We set the weight smoothing coefficient for the EMA of the teacher model to 0.9996. All experiments were executed on 4 NVIDIA A100 GPUs, with a batch size of 8 and a learning rate of 0.2.
@@
(b) Baseline Methods. We compared our results for 4 settings: (1) Lower Bound: supervised only training on the blended source data, (2) UDA: source domains are blended to apply state-of-art UDA methods, (3) MSDA: state-of-art MSDA methods, and (4) Upper Bound. We consider 3 upper bounds: Target Only: supervised training on target data, All-Combined: supervised training on target and source combined, Fine-Tuning: trained on source domain, then fine-tuned on the target domain.
@@
(c) Datasets and Adaptation Setting. For our experiments we use 5 different datasets with various objects and different domains and capturing conditions: BDD100K [37], Cityscapes [9], KITTI [12], MS COCO [19] and Synscapes [32]. The MSDA problem settings used in the empirical study are summarized in supplementary materials.
4.2 Results and Discussion
@@
(a) Cross-time Adaptation. Daytime and Night subsets of the BDD100K dataset are used as the source domain, while the Dusk/Dawn subset of BDD100K is used as the target domain. In this setting, the images are captured at different times. This causes a domain shift due to variations in illumination. The performance of ACIA is compared against other baselines in Tab. 1. Globally, UDA methods outperform the Lower Bound as expected, but there is a slight decrease in performance when the night domain is added. This is because the UDA Blending methods do not consider the domain shift among the source domains, resulting in performance degradation. This emphasizes the importance of considering MSDA methods. ACIA outperforms PMT [3] by 2.6 mAP. Also, our method outperforms Target-Only and All-Combined Upper Bound settings. The poor performance of the Target Only setting is due to fewer samples in the target domain compared to source domains.
Setting Method Source Data
D N D+N
Lower Bound Source Only 30.4 25.0 28.9
UDA Strong-Weak[27] 31.4 26.9 29.9
Graph Prototype[4] 31.8 27.6 30.6
Cat. Regularization[34] 31.2 28.4 30.2
UMT[10] 33.8 21.6 33.5
Adaptive Teacher[17] 34.9 27.8 34.6
MSDA MDAN[39] - - 27.6
M3SDA[22] - - 26.5
DMSN[36] - - 35.0
TRKP[33] - - 39.8
PMT[3] - - 45.3
ACIA (Ours) - - 47.9
Upper Bound Target-Only - - 26.6
All-Combined - - 45.6
Fine-Tuning - - 50.9
Table 1: mAP performance of ACIA and baseline methods for cross-time adaptation on BDD100k, when the 2 sources are daytime (D) and night (N) subsets, and the target is the Dusk/Dawn subset.
@@ D N D+N
(b) Cross-camera Adaptation. Cityscapes and Kitti datasets are used as source domains, while the Daytime subset of BDD100K is used as the target domain. In this setting, images are captured using different cameras, which causes a domain shift due to variations in capture conditions, like resolution and viewpoint. Training and evaluation are done only on the car category, as the only overlapping class across domains. The AP on the car compared against the baselines is reported in Tab. 2. The general variation of performance in this setting is similar to the Cross-time setting. Our proposed method outperforms the previous method by 0.4%. Our result is also very close to the target-only upper-bound. In this setting there is only one object category, thus the class conditioning has no impact on instance alignment.
Setting Method Source Data
C K C+K
Lower Bound Source Only 44.6 28.6 43.2
UDA Strong-Weak[27] 45.5 29.6 41.9
Cat. Regularization[34] 46.5 30.8 43.6
UMT[10] 47.5 35.4 47.0
Adaptive Teacher[17] 49.8 40.1 48.4
MSDA MDAN[39] - - 43.2
M3SDA[22] - - 44.1
DMSN[36] - - 49.2
TRKP[33] - - 58.4
PMT[3] - - 58.7
ACIA (Ours) - - 59.1
Upper Bound Target-Only - - 60.2
All-Combined - - 69.7
Fine-Tuning - - 72.1
@@ C K C+K
Table 2: AP for the car class of ACIA and baselines methods for the cross-camera setting. Source domains are Cityscapes (C) and Kitty (K) and the target is Daytime subset of BDD100K.
@@
(c) Extension to Mixed Domain Adaptation. The MS COCO, Cityscapes, and Synscapes datasets are used as source domains, while the Daytime subset of BDD100K is used as the target domain. In this setting, domain shift occurs due to multiple reasons. Results in Tab. 3 indicate that the Lower Bound outperforms UDA Blending baselines due to a large domain shift among domains. Thus, blending all the source domains as one and using UDA results in performance degradation. This highlights the difficulty of this setting. When Cityscapes and MS COCO are used as source domains, our method outperforms the previous state-of-the-art (PMT) by 2.3 mAP. When Synscapes is also added to the source domain this number rises to 2.6 mAP, indicating that our method improves even further when more source domains are added.
Setting Method Source Data
C C+M C+M+S
@@
Table 3: mAP performance on 7 object categories of ACIA and baseline methods in the mixed setting. Source domains are Cityscapes (C), MS COCO (M), and Synscapes (S) datasets while the Daytime domain of BDD100K is the target domain.
@@ 4.3 Ablation Studies
Unless mentioned, all ablation studies are performed on the cross-time adaptation setting as in [36, 3]. The impact of key components performing image-level and instance-level feature alignment is analyzed along with other possible design choices for class-conditioned instance alignment.
(a) Impact of the adaptation components. The impact of the individual components of our proposed method is analyzed – image-level discriminator and instance-level discriminator. To understand the importance of class conditioning in instance alignment, we examined the instance classifier with and without class-wise attention. Results are reported in Tab. 4. By adding the image-level discriminator, the performance improved considerably. Furthermore, by adding the instance-level discriminator (without class-embedding), the performance of the model increased by 3.4%, showing the importance of instance-level alignment. The performance is further increased by integrating the class information into the instance-level alignment.
@@
w/o class-embed w/ class-embed
@@
Table 4: Impact of individual ACIA components. Image-LA: Image-Level Alignment, Instance-LA: Instance-Level Alignment.
(b) Level of imbalance. In Tab. 5 compares ACIA against PMT for different levels of imbalance. When the dataset is balanced, both methods perform similarly. However, when the level of imbalance increases, ACIA obtains higher mAP. This shows that the error accumulation from the pseudo-labels is more problematic with imbalanced datasets. Our class-conditioned attention model mitigates this issue because it uses the same model for all classes, and therefore can better deal with high levels of imbalance.
Method Bike Bus Car Motor Person Rider Light Sign Truck mAP
Mostly Balanced 25.0% 50.0% 10.0% 15.0%
PMT 43.8 52.3 - 28.0 - 34.4 - - - 39.6
ACIA (Ours) 45.8 48.9 - 30.7 - 31.0 - - - 39.1
Slightly Imbalanced 4.6% 9.2% 1.8% 59.6% 2.8% 22.0%
PMT 42.1 52.7 - 25.1 43.1 34.0 - - 50.5 41.2
ACIA (Ours) 44.3 54.2 - 25.3 45.3 40.6 - - 50.4 43.4
Mostly Imbalanced 1.2% 2.4% 0.5% 15.4% 0.7% 32.8% 42.3% 5.7%
PMT 49.2 56.2 - 29.6 49.9 32.0 43.2 55.8 55.3 46.4
ACIA (Ours) 52.1 59.2 - 29.2 53.1 36.1 46.2 58.1 57.1 48.9
Highly Imbalanced 0.5% 1.0% 57.8% 0.2% 6.5% 0.3% 13.4% 17.8% 2.4%
PMT 55.3 59.8 67.6 29.9 47.6 32.7 46.3 56.0 57.7 50.3
ACIA (Ours) 56.1 61.0 69.2 31.9 51.8 39.8 49.2 59.0 61.0 53.2
Table 5: Detection results per class on the cross-time setting for PMT and our method with different levels of imbalance. The dataset is quite imbalanced. To create a more balanced dataset we reduced the number of classes by removing those that are creating the imbalance. For each level of imbalance, we report the relative importance of each class in percentage.
(c) Design choices for class-wise instance alignment, The class information can be incorporated into ROI-Pooled features in different ways, like concatenation and multiplication. Tab. 6 (left) reports the variation in performance with various design choices. Attention is the better design choice. We believe this is because just one class-embedding layer is learned across every domain. When merging the class information to the ROI-Pool features, attention can learn what information is more important for a particular domain in a better way than multiplication and concatenation. This is because the design of the attention layer allows it to learn complex relations [28].
@@
Table 6: (left) Different design choices for integrating class information in instance alignment; (right) Detection performance degrades when the target domain is used in instance alignment.
(d) Impact of target domain on class-wise alignment. In this section, we investigate the impact of employing or omitting the target image in instance-level alignment. As reported in Tab. 6 (right), utilizing target data in both scenarios leads to a decline in detector performance. In a DA problem, the target data lack labels, making it challenging to distinguish between foreground regions and object categories. One approach to address this challenge is by employing pseudo-labels, however, they often introduce noise and degrade performance. Nonetheless, in MSDA, the availability of multiple source domains enables us to choose which domains to align at the instance level. Hence, our work focuses on aligning solely the source domains.
(e) Detection visualization. In Fig. 4 we present three detection visualizations from no class-wise adaptation, PMT, and our ACIA. From the first two examples, it can be observed that our ACIA is better at detecting small objects that are harder to see. In the third example, our ACIA accurately detects a person, that has fewer samples in the dataset.
@@
(a) Methods w/o class alignment. (b) Prototype-based method. (c) Our ACIA method.
Figure 4: Examples of ODs on BDD100k cross-time setting, with different types of instance-level adaptation. (a) ODs with a multi-source adaptation without instance level class-conditional adaptation as in [36, 33]. (b) ODs with the prototype-based class-conditional adaptation [3]. (c) ODs with our attention-based class-conditional adaptation.